# A Vision Centric Remote Sensing Benchmark

### v2

@@ 2 Related Work
We explore the limitations of MLLMs in processing RS imagery. MLLMs leverage pre-trained LLMs [12, 3, 5] and integrate vision encoders [17, 24] to facilitate the alignment between visual and textual understanding. Language-visual alignment in such models can employ various techniques, such as simple projection layers[11], Q-Formers [8] and gated attention mechanisms [1]. Models like LLaVA and its variants [10, 11] highlight the role of high-quality instruction tuning data in enhancing multimodal learning for natural images. Unlike natural images, RS data exhibit fine-grained spatial structures in high resolution imagery, can include multiple modalities (e.g., hyperspectral data) and is impacted by environmental variations. Thus, the direct adaptation of MLLMs to RS data is challenging. Research on MLLMs for RS is largely based on LLaVA-based architectures, and several studies have used this framework to develop RS specific models, aiming to enhance vision-language alignment. GeoChat [6] and RS-LLaVA [2] instruction-tune LLaVA models on satellite imagery datasets, leveraging domain-specific ones such as RSVQA [14], FloodNet [18], DOTA [22] and RSIVQA-DOTA [26]. Besides these studies, EarthGPT [25] incorporates Synthetic Aperture Radar (SAR) and infrared modalities, addressing the challenges of multi-sensor fusion in RS tasks. Although, these models expand the scope of MLLMs for RS, yet their reliance on CLIP-based architectures limits their capacity to capture fine-grained spatial details and complex geospatial relationships.
@@ 3 Benchmark
A Visual Question Answering Benchmark for Remote Sensing MLLMs. To evaluate MLLMs’ ability to process RS imagery, we introduce a VQA benchmark specifically tailored for the identified CLIP-blind pairs. This benchmark enables us to assess how well MLLMs can visually ground information in high-resolution and complex RS imagery. The RSMMVP VQA dataset is constructed by analyzing 95 CLIP-blind image pairs, ensuring that it captures a diverse range of ambiguities and manually constructing the respective questions and choices. The benchmark consists of 300 questions within the VQA task, each designed to evaluate the models’ ability to distinguish fine-grained differences in RS images.
@@ 4 Experiments
Evaluation Protocols and Metrics. We adopt a paired question-answer assessment framework to ensure a rigorous evaluation of model accuracy. A model’s prediction was considered correct only if it accurately answered both questions associated with an image pair. This evaluation standard provides a robust measure of the ability of MLLMs to distinguish fine-grained visual differences in RS imagery. We use the 300 image-question pairs in our RSMMVP dataset and evaluate the accuracy of each model evaluated using GPT-4o that can assess the correctness of the answer with respect to the ground-truth answer.
@@
Benchmarking MLLMs on RSMMVP. To thoroughly evaluate the ability of MLLMs in processing RS images, we conduct a comprehensive benchmarking study on the RSMMVP dataset. The evaluation results in Figure 4 demonstrate a significant gap between human performance and the MLLMs’ performance. Human participants achieved an accuracy of 91.7%, which serves as an upper bound for the performance on our RSMMVP dataset. Among the models, GPT-4o (45.3%) achieves the highest accuracy, yet it shows significant performance gap with respect to the human performance. In contrast, the domain-adapted model RS-LLaVA-1.5-7B (26%) exhibits slight improvements over the general-purpose counter-part LLaVA-1.5-7B (22%), but still falls short in delivering robust performance for RS-specific applications. Notably, GeoChat-7B (16%) showed the lowest accuracy among the domain-adapted models. This result suggests that despite being fine-tuned on RS datasets, GeoChat lacks generalization capabilities. Finally, the LLaVA-1.5-13B achieved the second best result at 32%.

### v3

@@ 2 Related Work
We explore the limitations of MLLMs in processing RS imagery. MLLMs leverage pre-trained LLMs [12, 3, 5] and integrate vision encoders [17, 24] to facilitate the alignment between visual and textual understanding. Language-visual alignment in such models can employ various techniques, such as simple projection layers[11], Q-Formers [8] and gated attention mechanisms [1]. Models like LLaVA and its variants [10, 11] highlight the role of high-quality instruction tuning data in enhancing multimodal learning for natural images. Unlike natural images, RS data exhibit fine-grained spatial structures in high resolution imagery, can include multiple modalities (e.g., hyperspectral data) and is impacted by environmental variations. Thus, the direct adaptation of MLLMs to RS data is challenging. Research on MLLMs for RS is largely based on LLaVA-based architectures, and several studies have used this framework to develop RS specific models, aiming to enhance vision-language alignment. GeoChat [6] and RS-LLaVA [2] instruction-tune LLaVA models on satellite imagery datasets, leveraging domain-specific ones such as RSVQA [14], FloodNet [18], DOTA [22] and RSIVQA-DOTA [26]. Besides these studies, EarthGPT [25] incorporates Synthetic Aperture Radar (SAR) and infrared modalities, addressing the challenges of multi-sensor fusion in RS tasks. Although, GeoChat and RS-LlaVA models expand the scope of MLLMs for RS, yet their reliance on CLIP-based architectures limits their capacity to capture fine-grained spatial details and complex geospatial relationships.
@@ 3 Benchmark
A Visual Question Answering Benchmark for Remote Sensing MLLMs. To evaluate MLLMs’ ability to process RS imagery, we introduce a VQA benchmark specifically tailored for the identified CLIP-blind pairs. This benchmark serves as a proxy evaluation of MLLMs’ visual grounding abilities in high-resolution and complex RS imagery. The RSMMVP VQA dataset is constructed by analyzing 95 CLIP-blind image pairs, ensuring that it captures a diverse range of ambiguities and manually constructing the respective questions and choices. The benchmark consists of 300 questions within the VQA task, each designed to evaluate the models’ ability to distinguish fine-grained differences in RS images.
@@ 4 Experiments
Evaluation Protocols and Metrics. We adopt a paired question-answer assessment framework to ensure a rigorous evaluation of model accuracy. A model’s prediction was considered correct only if it accurately answered both questions associated with an image pair. This evaluation standard provides a robust measure of the ability of MLLMs to distinguish fine-grained visual differences in RS imagery. We use the 300 image-question pairs in our RSMMVP dataset and evaluate the accuracy of each model using GPT-4o.
@@
Benchmarking MLLMs on RSMMVP. To thoroughly evaluate the ability of MLLMs in processing RS images, we conduct a comprehensive benchmarking study on the RSMMVP dataset. The evaluation results in Figure 4 demonstrate a significant gap between human performance and the MLLMs’ performance. Human participants achieved an accuracy of 91.7%, which serves as an upper bound for the performance on our RSMMVP dataset. Among the models, GPT-4o (45.3%) achieves the highest accuracy, yet it shows significant performance gap with respect to the human performance. In contrast, the domain-adapted model RS-LLaVA-1.5-7B (26%) exhibits slight improvements over the general-purpose counter-part LLaVA-1.5-7B (22%), but still falls short in delivering robust performance for RS-specific applications. Notably, GeoChat-7B (16%) showed the lowest accuracy among the domain-adapted models. Note that the reported result for GeoChat corresponds to a re-trained version of the model, excluding the images used in constructing the benchmark. This result suggests that despite being fine-tuned on RS datasets, GeoChat lacks generalization capabilities. Finally, the LLaVA-1.5-13B achieved the second best result at 32%.